OpenStudy (anonymous):

MEDAL AND FAN PLEASE HELP DESPERATE!!! Simplify cube root of 5 over fourth root of 5.

OpenStudy (anonymous):

$\mathrm{Simplify}\:3\cdot\sqrt{5}\cdot\sqrt{5}:\quad15$Simplify 3·√5·√5: 15 $=\frac{15}{4}$=154

OpenStudy (anonymous):

=15\4 is correct

OpenStudy (agent0smith):

Change the roots into fractional exponents. Like cube root becomes a1/3 exponent. Then use exponent rules to simplify.

OpenStudy (anonymous):

That's not an answer choice ^ and can you elaborate, agent?

OpenStudy (agent0smith):

Cube root is the same thing as a 1/3 exponent. Fourth root is the same as a1/4 exponent. So do that first.

OpenStudy (anonymous):

The answer choices are: 5 1/4, 5 1/12, 5 7/12, 5 4/3

OpenStudy (anonymous):

how do I do that?

OpenStudy (agent0smith):

Cube root of 5 becomes 5^1/3. Do the same with fourth root of 5.

OpenStudy (anonymous):

So the first option?

OpenStudy (agent0smith):

I'm giving you step one of like three or four steps.

OpenStudy (agent0smith):

Cube root of 5 is the same thing as 5^1/3. You tell me: the fourth root of 5 is the same as...

OpenStudy (anonymous):

I'm sorry dude, Im just kinda rushing cause I have very limited time :/

OpenStudy (mathmale):

Confused: If you truly want to learn this material, you'll have to set aside enough time to learn it. Kindly do not try to rush people who have been trying to help you, and do not ask for or expect to receive direct answers.

OpenStudy (mathmale):

Simplify:

OpenStudy (mathmale):

\[\sqrt[3]{5}\div \sqrt[4]{5}\]

OpenStudy (mathmale):

Agent0Smith has already given you the key info you need to solve this problem. Might want to go back and re-read what he has advised.